OpenStudy (anonymous):

Mario has 5 times his son's age. Three years ago Mario was 5 times his son's age. What are their ages? Help me please

OpenStudy (anonymous):

x= 3/5

OpenStudy (anonymous):

Let: x = marios age now y = son's age So... x = 5y (x-3) = 5(y-3) Plugging that in: 5y - 3 = 5y - 15 Something's wrong...?

OpenStudy (anonymous):

since mario is 5 times something = 5x and when they say 3 years ago so it's -3 = 5x - 3 = 0, solve for x and x = 3/5 I guess :)

OpenStudy (anonymous):

hmm, question is confusing, there's something wrong indeed.

OpenStudy (anonymous):

And you're supposed to realize that before you put in an answer, you know. Do the work, and get it correct, like you said. Hypocrite!!!

OpenStudy (anonymous):

check if something is wrong with the question....at both times Mario is 5 X son's age?

OpenStudy (anonymous):

Sorry. This is the correct problem Mario has 4 times his son's age. Three years ago Mario was 5 times his son's age. What are their ages? Help me please

OpenStudy (anonymous):

solve x = 4y and (x-3) = 5(y-3). you'll get x=48, y=12

OpenStudy (anonymous):

Thank you guys, I finally solved it.

OpenStudy (anonymous):

Yeah, not cool giving out the wrong question.

OpenStudy (anonymous):

Sorry guys. You must understand